Two sides and an angles are given. Determine whether the given information results in one triangle, two triangles , or no triangle at all. If there is one or more triangles solve any triangle(s) that results. If there is no triangle, show and provide an explanation why.


B = 106 degrees, b = 5, a = 23


1
Expert's answer
2021-05-31T06:58:41-0400

Ans:- Given "B = 106 ^{\circ} ,\ b = 5, \ a = 23"

We have using the Law of Sines....

"\Rightarrow \dfrac{SinA}{a}=\dfrac{SinB}{b}"


"\Rightarrow \dfrac{SinA}{23}=\dfrac{Sin106^{\circ}}{5}"


So.........

"Sin^{-1}(23\times\dfrac{Sin106^{\circ}}{5})=A=" Undefined, outside the range of "Sin\ function" .

Therefore Angle A does not exist.


So There is no triangle exist of that type having parameters "B = 106 ^{\circ}, \ b = 5, \ a = 23"